# PriorDiffusion: Leverage Language Prior in Diffusion Models

### v1

@@ Abstract
This paper explores the potential of leveraging language priors learned by text-to-image diffusion models to address ambiguity and visual nuisance in monocular depth estimation. Particularly, traditional monocular depth estimation suffers from inherent ambiguity due to the absence of stereo or multi-view depth cues, and nuisance due to lack of robustness of vision. We argue that language prior in diffusion models can enhance monocular depth estimation by leveraging the geometric prior aligned with the language description, which is learned during text-to-image pre-training. To generate images that reflect the text properly, the model must comprehend the size and shape of specified objects, their spatial relationship, and the scale of the scene. Thus, we propose PriorDiffusion, using a pre-trained text-to-image diffusion model that takes both image and text description that aligned with the scene to infer affine-invariant depth through a denoising process. We also show that language priors can guide the model’s attention to specific regions and help it perceive the 3D scene in alignment with user intent. Simultaneously, it acts as a constraint to accelerate the convergence of the diffusion trajectory, since learning 3D properties from a condensed, low-dimensional language feature is more efficient compared with learning from a redundant, high-dimensional image feature. By training on HyperSim and Virtual KITTI, we achieve state-of-the-art zero-shot performance and a faster convergence speed, compared with other diffusion-based depth estimators, across NYUv2, KITTI, ETH3D, and ScanNet.
@@ 1 Introduction
Monocular depth estimation requires the model to predict pixel-wise depth from a single image. Recent advancements in diffusion-based models have shown promise for generating high-quality depth maps [12, 34, 66, 67, 109, 35] due to their ability to capture complex structures and fine details by progressively denoising a latent representation to produce accurate depth predictions. However, it remains a challenging task due to inherent ambiguities and various visual nuisances. One of the many inherent ambiguities is the scale ambiguity. Unlike stereo or multi-view depth estimation methods, monocular approaches lack the geometric depth cues provided by multiple perspectives, leading to difficulty in accurately discerning the relative sizes and distances of objects within a scene. That is, different real-world 3D scenes can project to the same 2D image under different scales. For example, a small object up close can appear identical to a larger object far away in an image. Another challenging ambiguity is texture ambiguity. Surfaces that are uniformly textured or have repetitive patterns, like tiled floors, can create confusion for depth estimators. The models might misinterpret the repetition as uniform depth, even though the actual depth might vary (e.g., a tiled floor receding into the distance). Furthermore, visibility ambiguity creates limitations for discerning small, partial or occluded objects purely from the visual signal. On the other hand, visual nuisances such as illumination, motion blur, atmospheric conditions, and perspective distortion distort or obscure critical visual cues, leading vision algorithms to misinterpret scene structure, depth, and object boundaries, thus reducing their accuracy and reliability.
@@
An emerging solution lies in leveraging language as a complementary modality to provide a prior to resolve ambiguity and visual nuisance. Language descriptions provided by humans offer a wealth of priors about spatial relationships, scene types, object sizes, and depth hierarchies. For example, distinguishing between a nearby toy car and a distant real car can be challenging for a model. However, when human-provided language inputs specify details like “This is a close toy car,” this serves as a prior that informs the model about the 3D size and shape of the object. Combined with the observed size, this prior helps narrow down the resolution space for depth estimation, leading to more accurate depth predictions. In cases where an object is small or partially occluded, language can indicate the object’s identity and semantics, enabling the model to infer its typical 3D size and shape and make more precise depth predictions for the observed portion. Language inputs are also valuable in resolving texture ambiguities. Descriptive cues can specify the type and function of surfaces in a scene, aiding the model in inferring geometric properties. For instance, if a uniformly textured surface is described as a “white ceiling,” the model can interpret it as a continuous, flat plane at a consistent depth above the observer, extending into the distance. Similarly, for surfaces with repetitive patterns, such as a red and black interwoven carpet, language inputs like “A red and black interwoven carpet extending across the room” can help the model recognize that the repeating pattern belongs to a flat surface, preventing misinterpretation of edges between tiles as depth discontinuities.
@@
To this end, we propose PriorDiffusion, incorporating human-provided language descriptions of the scene as priors to guide the depth map predication. During text-to-image pre-training, diffusion models learn to generate diverse scenes that align with the provided language descriptions. To achieve this, the diffusion model must comprehend the size and shape of each object, their spatial relationship, and the scale of the scene described in the language input, to generate images that accurately represent the input text under various viewpoints and scene layouts. Thus, for depth estimation, the language priors provided by human-generated captions allow the model to leverage the geometric priors embedded in the input text to more effectively perceive the 3D scene. In our PriorDiffusion, during the denoising process, both the image and the language input are used by the model to predict the noise to be removed. In the end, Gaussian noise is progressively refined into a depth map that aligns with both the input image and the language description. Given the difficulty of obtaining sufficient human-provided text descriptions for every training and inference image, we utilize a vision-language model, i.e. LLaVA [11], to generate descriptions for each image, simulating human annotation.
To support our hypothesis, we conduct training on two synthetic datasets, HyperSim [62] and Virtual KITTI [4], and conduct zero-shot evaluation on four real-world datasets, NYUv2 [69], KITTI [24], ETH3D [68], and ScanNet [10]. Comparing with other state-of-the-art diffusion depth estimators, we achieve superior qualitative results, and higher-fidelity quantitative results, with faster convergence speeds.
@@
Our contributions. By leveraging language as a prior for monocular depth estimation using a diffusion model pretrained with text-to-image generation, we demonstrate that:
@@
  • Language can provide a prior about the existence, geometric properties, and spatial relationships of objects specified in the language description, allowing depth estimators to better infer the depth of those objects.
@@
  • Language priors direct the model’s attention to specific regions, guiding it to perceive the 3D scene according to the user’s intent. This is particularly beneficial for regions that pose challenges to vision systems, such as those with small size, poor illumination, occlusion, or high visual similarity to the background.
  • Language priors serve as a constraint to speed up the convergence of diffusion trajectory. Our intuition is that learning 3D properties like geometric characteristics and spatial relationships from a condensed, low-dimensional language feature is more efficient compared with learning from redundant, high-dimensional image features.
@@ 2 Related Work
Monocular depth estimation. Monocular depth estimation is a dense regression problem, requiring pixel-wise predictions of depth from a single image, capturing complex spatial relationships within the scene. Some depth models learn to infer pixel-wise depth in metric scale (i.e. meters) [2, 19, 36, 46, 80, 95, 105, 96] by minimizing loss between depth predictions and ground-truth depth maps. Each model typically applies to only one data domain in which it is trained, with similar camera parameters and object scales. Specifically, DORN [19] leverages a spacing-increasing discretization technique. AdaBins [2] partitions depth ranges into adaptive bins. NeWCRFs [95] uses neural window fully-connected CRFs to compute energy. VA-DepthNet [46] uses first-order variational constraints, making the network aware of the depth gradient in the scene space. When ground-truth depth is not available, self-supervised approaches [3, 18, 33, 37, 38, 51, 55, 74, 76, 77, 79, 84, 94, 100, 106, 107, 111, 114] rely on geometric constraints, where scale is attributed through radar [70], inertials [28, 78], lidar [16, 49, 54, 83, 82, 81, 85, 91], or binocular images [23, 27, 26]. To enable generalization across diverse domains, empowering depth estimation in the wild, other works have explored affine-invariant depth estimation [98, 88, 60, 59, 35, 89, 93, 13, 92, 45], which predicts depth up to an unknown global shift and scale. This approach can accommodate varied scenes with different scales, while still preserving the geometric relationships between different objects or regions within the scene. HND [98] hierarchically normalizes the depth representations with spatial information and depth distributions. Depth Anything  [88] learns from large-scale automatically annotated data. DPT [60] leverages vision transformers using a scale- and shift-invariant trimmed loss. MiDas [59] mixes multiple datasets with training objectives invariant to depth range and scale. Marigold [35] associates fine-tuning protocol with a diffusion model. Existing methods try to predict depth solely from images, which suffers from ambiguity like scale, object orientation, occlusion, and visual nuisance like viewpoints, illumination, appearance, texture, etc. On the other hand, our PriorDiffusion leverages language prior in Diffusion Models to provide a thorough understanding of the scene associated with the input text, to resolve the ambiguity and the nuisance variables in the visual algorithm.
@@
Diffusion model for monocular depth estimation. Denoising Diffusion Probabilistic Model (DDPM) [30] learns to reverse a diffusion process that progressively degrades images with Gaussian noise so that they can draw samples from the data distribution by applying the reverse process to random noise. They have been applied to various tasks like text-to-image geneartion [63, 58, 118, 117, 75], super resolution [64, 41, 22], image inpainting [50, 9, 90], 3D object generation [56, 40, 48], etc. Several approaches have explored the use of DDPM for metric depth estimation. DiffusionDepth [12] learns to denoise random depth distribution into a depth map with monocular visual conditions. Similarly, DDP [34] encodes the input image and use diffusion to decodes it into a depth map. DepthGen [66] extends a multi-task diffusion model for metric depth prediction. Its successor, DDVM [67], focuses on pretraining with synthetic and real datasets to improve depth estimation performance. Lastly, VPD [109] and Marigold [35] utilizes a pretrained Stable Diffusion model [63] as an image feature extractor, incorporating additional text input for enhanced depth prediction. While previous works only focused on using images as conditions of diffusion models to predict depth, our PriorDiffusion utilizes language as priors in diffusion models to resolve ambiguity and visual nuisances, and conduct 3D perception in a controllable way.
@@
Language for monocular depth estimation. Vision-Language models  [5, 43, 42, 53, 57, 73] acquire a comprehensive understanding of languages and images through pre-training under diverse datasets, thereby establishing a robust foundation for downstream tasks. [86, 87, 103, 7, 72, 17, 104, 105, 116, 96, 97, 44, 11].
Typically, CLIP [57] employs contrastive learning on text-image pairs, enabling a range of applications like few-shot image classification  [21, 101, 102, 113], image segmentation  [61, 110], object detection  [61, 115], and 3D perception [32, 104, 105, 116, 112, 99, 8, 29]. In light of their emerging ability, some works [1, 32, 105, 108] have tried to apply vision-language models for monocular depth estimation. Particularly, DepthCLIP [105] leverages the semantic depth response of CLIP with a depth projection scheme to conduct zero-shot monocular depth estimation. WorDepth [96] learned the variational prior distribution of 3D scenes from text captions. RSA [97] predict scale using language description to align relative depth to metric scale. While language has been applied in various aspects of 3D perception, our work is the first to use language priors in diffusion models to leverage the geometric prior aligned with the language description to enhance 3D perception.
@@ 3 Method
Diffusion formulation. To use a diffusion model for monocular depth estimation, following [12, 34, 66, 67, 109, 35], we adopt the formulation of Denoising Diffusion Probabilistic Model (DDPM) [30]. We define a latent variable $\mathbf{z}_{t}$ $\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times C^{\prime}}$that represents a noisy version of the depth map in the latent space at diffusion timestep $t$. The ground-truch depth map $y^{*}$ is encoded into a latent representation $\mathbf{z_{y}}$ as supervision using a frozen VAE encoder $E$: $\mathbf{z_{y}}=E(y^{*})$. The diffusion latent after the whole denoising process $\mathbf{z_{0}}$ is decoded back to the image space using the frozen VAE decoder $D$: $\hat{y}=D(\mathbf{z}_{0})$.
Figure 4: Visualization on KITTI. Compared to Margold, our PriorDiffusion model shows superior accuracy in depth prediction for a given input image, especially for objects highlighted in the language description (marked in red), even when parts of the object are almost visible in the image (such as the parking car in the first column). This shows that the language prior can help the model better infer the geometric properties of these objects and improve distance estimation. Furthermore, the language descriptions guide the model’s attention to specific regions relevant to user needs, such as a sign at a distance, which can potentially enhance the safety of self-driving systems that rely solely on vision sensors.
@@
where $\mu_{\theta}$ and $\Sigma_{\theta}$ is the diffusion model, that predicts the mean and variance conditioned on $\mathbf{z}_{t}$, $t$, the input image $x$ and the corresponding text $\mathbf{c}$. We use the U-Net from Stable Diffusion v2 [63] to initialize is the diffusion model. Specifically, given a text caption $\textbf{c}=\{c_{1},c_{2},...\}$, we first encode it through a frozen CLIP text encoder, then feed it into the diffusion model. Given an image $x$, we encode it using the same VAE encoder $E(x)$ that encoded the depth map, concatenated with the depth latent $\mathbf{z}_{t}$, then feed it into the diffusion model.
@@
Method NYUv2 KITTI ETH3D ScanNet $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ $\delta_{1}$ AbsRel↓ DiverseDepth [92] 87.5 11.7 70.4 19.0 69.4 22.8 88.2 10.9 MiDaS [59] 88.5 11.1 63.0 23.6 75.2 18.4 84.6 12.1 LeReS [93] 91.6 9.0 78.4 14.9 77.7 17.1 91.7 9.1 Omnidata [13] 94.5 7.4 83.5 14.9 77.8 16.6 93.6 7.5 HDN [98] 94.8 6.9 86.7 11.5 83.3 12.1 93.9 8.0 DPT [60] 90.3 9.8 90.1 10.0 94.6 7.8 93.4 8.2 Marigold [35] 95.9 6.0 90.4 10.5 95.1 7.1 94.5 6.9 Marigold (iter: 20000) 95.2 6.4 86.2 12.0 93.9 7.6 94.8 7.1 Marigold (iter: 30000) 95.7 6.1 89.7 10.7 95.4 6.9 94.0 7.3 Ours (iter: 20000) 95.9 6.0 90.4 10.4 95.6 6.6 94.4 7.1 Ours (iter: 30000) 95.9 5.9 90.6 10.4 95.7 6.5 94.9 6.8
Table 1: Quantitative Comparison. We compare our PriorDiffusion against other state-of-the-art affine-invariant depth estimators across various zero-shot benchmarks. Metrics are reported as percentages. * indicates the results of the model that we re-implemented. Note that our method converges faster compared with Marigold, and is even better with only two-thirds iterations, attributing the language prior which provides a stronger constraint for the fast convergence of the diffusion trajectory.
@@
Obtain language description. To implement our method, we require human-provided text descriptions for each image. Since standard benchmarks do not include such descriptions, we propose using off-the-shelf models to generate text descriptions to simulate those a human would provide. To closely mimic human-like descriptions, we focus on natural text that is not constrained by templates and better resembles human input. For this purpose, we use the visual question-answering model LLaVA v1.6 Mistral [47]. We prompt this model with:
This prompt is designed to elicit responses that include essential details, such as the positioning of objects, their interactions, and notable features that may impact depth estimation. Note that all training data we use are synthetic data, and by emphasizing “assuming it’s a real-world image,” we ensure that the descriptions align with the types of inputs and scenarios the model will encounter.

### v2

@@ Abstract
Traditional monocular depth estimation suffers from inherent ambiguity and visual nuisance. We argue that language prior can enhance monocular depth estimation by leveraging the inductive bias learned during the text-to-image pre-training of diffusion models. The ability of these models to generate images that align with text indicates that they have learned the spatial relationships, size, and shape of specified objects, which can be applied to improve depth estimation. Thus, we propose PriorDiffusion, using a pre-trained text-to-image diffusion model that takes both images and corresponding text descriptions to infer affine-invariant depth through a denoising process. We also show that language prior enhances the model’s perception of specific regions of images that users care about and describe. Simultaneously, language prior acts as a constraint to accelerate the convergence of both training and the inference diffusion trajectory. By training on HyperSim and Virtual KITTI, we achieve faster training convergence, fewer inference diffusion steps, and state-of-the-art zero-shot performance across NYUv2, KITTI, ETH3D, and ScanNet. Code will be released upon acceptance.
@@ 1 Introduction
Monocular depth estimation requires the model to predict pixel-wise depth from a single image. Recent advancements in diffusion-based models have shown promise for generating high-quality depth maps [12, 37, 69, 70, 110, 38, 19, 29, 17, 23, 31, 87] due to their ability to capture complex structures and fine details by progressively denoising a latent representation to produce accurate depth predictions. However, it remains a challenging task due to inherent ambiguities and various visual nuisances. Unlike stereo or multi-view depth estimation methods, monocular approaches lack depth cues provided by multiple perspectives, leading to difficulty in accurately discerning the relative sizes, distances, and spatial relationships of objects within a scene. That is, different real-world 3D scenes can project to the same 2D image, causing inherent ambiguities. For example, texture ambiguity. Surfaces that are uniformly textured or have repetitive patterns, like tiled floors, can create confusion for depth estimators. The models might misinterpret the repetition as uniform depth, even though the actual depth might vary (e.g., a tiled floor receding into the distance). Also, visibility ambiguity creates limitations for discerning small, partial or occluded objects purely from the visual signal. On the other hand, visual nuisances such as illumination, motion blur, atmospheric conditions, and perspective distortion distort or obscure critical visual cues, leading vision algorithms to misinterpret scene structure, depth, and object boundaries, thus reducing their accuracy and reliability.
Shown in Figure 1, an emerging solution lies in leveraging language as a complementary modality to provide a prior to resolve ambiguity and visual nuisance. Language descriptions provided by humans offer a wealth of priors about spatial relationships, scene types, object sizes, and depth hierarchies.
@@
In cases where an object is small, partially occluded, visually similar to the background, or its type is ambiguous from a single image, language can indicate the object’s identity and semantics, enabling the model to infer its typical 3D size and shape and make more precise depth predictions for the observed portion. Language inputs are also valuable in resolving texture ambiguities. Descriptive cues can specify the type and function of surfaces in a scene, aiding the model in inferring geometric properties. For instance, if a uniformly textured surface is described as a “white ceiling”, the model can interpret it as a continuous, flat plane at a consistent depth above the observer, extending into the distance. Similarly, for surfaces with repetitive patterns, such as a red and black interwoven carpet, language inputs like “A red and black interwoven carpet extending across the room” can help the model recognize that the repeating pattern belongs to a flat surface, preventing misinterpretation of edges between tiles as depth discontinuities. As shown in Figure 2, better perception can be achieved for insignificant regions indicated in the language description.
@@
To this end, we propose PriorDiffusion, incorporating human-provided language prior of the scene as an inductive bias to guide the depth map prediction. During text-to-image pre-training, diffusion models learn to generate diverse images under various viewpoints and scene layouts that align with the provided language descriptions. The ability of these models to generate images that align with text suggests they have learned the spatial relationships, size, and shape of specified objects in the described 3D scene, which can be leveraged to improve depth estimation. In our PriorDiffusion, during the denoising process, both the image and the text input are used by the model as conditions to predict the noise to be removed. Ultimately, Gaussian noise is progressively refined into a depth map that aligns with both the input image and the language description. In real-world applications when embodied agents are integrated with PriorDiffusion, humans can describe the environment to help agents better perceive it. When generalizing to a new environment, language, having less covariate shift compared to images, allows humans to describe the new environment, aiding agents in better adapting and generalizing to it.
To support our hypothesis, we conduct training on two synthetic datasets, HyperSim [65] and Virtual KITTI [4], and conduct zero-shot evaluation on four real-world datasets, NYUv2 [72], KITTI [25], ETH3D [71], and ScanNet [10]. Given the difficulty of obtaining sufficient human-provided text descriptions for every training and inference image, we utilize a vision-language model, i.e. LLaVA [11], to generate descriptions for each image, simulating human annotation. we achieve faster training convergence, fewer inference diffusion steps, and state-of-the-art zero-shot performance across NYUv2, KITTI, ETH3D, and ScanNet.
@@
Our contributions. By leveraging language prior for diffusion-based monocular depth estimation, we demonstrate that:
@@
  • Language prior provides an inductive bias about the existence, geometric properties, and spatial relationships of objects specified in the language description, allowing depth estimators to better perceive the depth of those objects.
@@
  • Language prior directs the model’s attention to specific regions, guiding it to perceive the 3D scene according to the user’s intent. This is particularly beneficial for regions that pose challenges to vision systems, such as those with small size, poor illumination, occlusion, or high visual similarity to the background.
  • Language prior serves as a constraint to speed up training convergence, and serves as a good initialization of diffusion trajectory to speed up inference. This improves the efficiency of the diffusion model, optimizing both its performance and computational resource usage.
Figure 2: 3D reconstruction. For insignificant regions indicated in language description, PriorDiffusion achieves better depth estimates and reconstructions thanks to the use of language.
@@ 2 Related Work
Monocular depth estimation. Monocular depth estimation is a dense regression problem, requiring pixel-wise predictions of depth from a single image, capturing complex spatial relationships within the scene. Some depth models learn to infer pixel-wise depth in metric scale (i.e. meters) [2, 18, 39, 48, 81, 97, 106, 98] by minimizing loss between depth predictions and ground-truth depth maps. Each model typically applies to only one data domain in which it is trained, with similar camera parameters and object scales.
@@
When ground-truth depth is not available, self-supervised approaches [3, 17, 36, 40, 41, 53, 57, 75, 77, 78, 80, 85, 96, 102, 107, 108, 112, 115] rely on geometric constraints, where scale is attributed through radar [73], inertials [30, 79], lidar [16, 51, 56, 84, 83, 82, 86, 93], or binocular images [24, 28, 27]. To enable generalization across diverse domains, empowering depth estimation in the wild, other works have explored affine-invariant depth estimation [100, 90, 63, 62, 38, 91, 95, 13, 94, 47], which predicts depth up to an unknown global shift and scale. This approach can accommodate varied scenes with different scales, while still preserving the geometric relationships between different objects or regions within the scene. HND [100] hierarchically normalizes the depth representations with spatial information and depth distributions. Depth Anything  [90] learns from large-scale automatically annotated data. DPT [63] leverages vision transformers using a scale- and shift-invariant trimmed loss. MiDas [62] mixes multiple datasets with training objectives invariant to depth range and scale. Depth Anything [90] and Depth Anything v2 [91] train a depth foundation model using large-scale unlabeled data. Existing methods try to predict depth solely from images, which suffers from ambiguity like scale, object orientation, occlusion, and visual nuisance like viewpoints, illumination, appearance, texture, etc. On the other hand, our PriorDiffusion leverages language prior in Diffusion Models to provide a thorough understanding of the scene associated with the input text, to resolve the ambiguity and the nuisance variables in the visual algorithm.
@@
Diffusion model for monocular depth estimation. Denoising Diffusion Probabilistic Model (DDPM) [33] learns to reverse a diffusion process that progressively degrades images with Gaussian noise so that they can draw samples from the data distribution by applying the reverse process to random noise. They have been applied to various tasks like text-to-image generation [66, 61, 119, 118, 76], super resolution [67, 44, 22], image inpainting [52, 9, 92], 3D object generation [58, 43, 50], etc. Several approaches have explored the use of DDPM for metric depth estimation. DiffusionDepth [12] learns to denoise random depth distribution into a depth map with monocular visual conditions. Similarly, DDP [37] encodes the input image and use diffusion to decodes it into a depth map. DepthGen [69] extends a multi-task diffusion model for metric depth prediction. Its successor, DDVM [70], focuses on pretraining with synthetic and real datasets to improve depth estimation performance. Build upon the diffusion model, DepthFM [29] uses flow matching to conduct a faster monocular depth estimation. In light of strong representation learned by pre-trained, text-to-image diffusion models, recent methods like VPD [110], E2E FT [23], GeoWizard [19], Lotus [31], GenPercept [87], and Marigold [38] utilize a pretrained Stable Diffusion model [66] as a backbone to predict depth through a denoising process. Specifically, Marigold [38] concatenates input images with diffusion latent as conditions for denoising U-Net to predict noise that needs to be removed, to gradually denoise a Gaussian noise to a depth map. While previous works only focused on using images as conditions of diffusion models to predict depth, our PriorDiffusion utilizes language prior as inductive bias for diffusion models to resolve ambiguity and visual nuisances, and conduct 3D perception in a controllable way.
@@
Language for monocular depth estimation. Vision-Language models  [5, 46, 45, 55, 60] acquire a comprehensive understanding of languages and images through pre-training under diverse datasets, thereby establishing a robust foundation for downstream tasks. [88, 89, 104, 7, 105, 106, 117, 98, 99, 11]. Typically, CLIP [60] employs contrastive learning on text-image pairs, enabling a range of applications like few-shot image classification  [21, 103, 59, 114], image segmentation  [64, 111], object detection  [64, 116], and 3D perception [35, 105, 106, 117, 113, 101, 8, 32]. In light of their emerging ability, some works [1, 35, 106, 109] have tried to apply vision-language models for monocular depth estimation. Particularly, DepthCLIP [106] leverages the semantic depth response of CLIP with a depth projection scheme to conduct zero-shot monocular depth estimation. WorDepth [98] learns the variational prior distribution of 3D scenes from text. RSA [99] predicts scale using language description to align relative depth to metric scale. While language has been applied in various aspects of 3D perception, our work is the first to use language prior for diffusion models to enhance 3D perception.
@@ 3 Method
Diffusion formulation. The pipeline of PriorDiffusion is shown in Figure 3. To use a diffusion model for monocular depth estimation, following [12, 37, 69, 70, 110, 38], we adopt the formulation of Denoising Diffusion Probabilistic Model (DDPM) [33]. We define a latent variable $\mathbf{z}_{t}$ $\in\mathbb{R}^{H^{\prime}\times W^{\prime}\times C^{\prime}}$that represents a noisy version of the depth map in the latent space at diffusion timestep $t$. The ground truch depth map $y^{*}$ is encoded into a latent representation $\mathbf{z_{y}}$ as supervision using a frozen VAE encoder $E$: $\mathbf{z_{y}}=E(y^{*})$. The diffusion latent after the whole denoising process $\mathbf{z_{0}}$ is decoded back to the image space using the frozen VAE decoder $D$: $\hat{y}=D(\mathbf{z}_{0})$.
Figure 5: Visualization on KITTI. PriorDiffusion predicts better depth for described objects, even when parts of the object are almost visible in the image (such as the parking car in the first column). The model is guided for specific described regions, such as a sign at a distance, potentially enhancing the safety of self-driving systems that rely solely on vision sensors.
@@
where $\mu_{\theta}$ and $\Sigma_{\theta}$ is the diffusion model, that predicts the mean and variance conditioned on $\mathbf{z}_{t}$, $t$, the input image $x$ and the corresponding text $\mathbf{c}$. We use the denoising U-Net from Stable Diffusion v2 [66] to initialize is the diffusion model. Specifically, given a text caption $\textbf{c}=\{c_{1},c_{2},...\}$, we first encode it through a frozen CLIP text encoder, then feed it into the diffusion model. Given an image $x$, we encode it using the same VAE encoder $E(x)$ that encoded the depth map, concatenated with the depth latent $\mathbf{z}_{t}$, then feed it into the diffusion model.
@@
[…]
Table 1: Quantitative comparison. PriorDiffusion achieves performance on par with other state-of-the-art affine-invariant depth estimators that use significantly more training images, and outperforms methods trained with a similar number of images. Metrics are reported as percentages. * indicates the results of the model that we re-implemented. Note that our method converges faster compared with Marigold, and is even better with only two-thirds iterations, attributing the language prior which provides a stronger constraint for the fast convergence of the diffusion trajectory. For fair comparison, PriorDiffusion and Marigold are evaluated without ensemble.